# Multimodal Adversarial Defense for Vision-Language Models by Leveraging One-To-Many Relationships

### v2

@@ 6.1.1 Effectiveness of Multimodal Adversarial Training
6.1.2 Effectiveness of one-to-many augmentations
We observe that effective augmentations, such as I2T(Human), consistently enhance robustness, whereas some augmentations, such as T2I(SD), do not.
To identify key factors for effective augmentation, we analyze alignment, diversity, and distribution gap based on our hypothesis for approximating the true image-text distribution (Sec.4.3.2). Figure 2 presents these scores, computed in CLIP’s embedding space. Alignment measures the cosine similarity between augmented image-text pairs, diversity quantifies the average KL-divergence between each original and augmented sample, and distribution gap measures the overall distance between the distributions of original and augmented samples. Figure 2 reveals that all properties, high alignment, high diversity, and small distribution gap, is crucial for effective augmentations.
Cross-modal augmentation outperforms intra-modal augmentation due to higher image-text alignment. We observe that cross-modal augmentations can result in better robustness than intra-modal augmentations. This can be attributed to cross-modal augmentations taking into account the other modality, generating well-aligned image-text pairs. For example, the alignment scores of RandAug and EDA are lower than those of the original pairs (Fig. 2), indicating semantic mismatch (Case. 1). Intra-modal augmentations struggle to balance diversity and alignment: increasing diversity without considering image-text relationships disrupts alignment. This is evident in Fig. 2, where intra-modal augmentations appear in the lower-left, achieving only one of alignment or diversity.
@@ 6.1.2 Effectiveness of one-to-many augmentations
Appendix B provides qualitative examples of the augmentations generated for each technique.
@@
Image augmentations are challenging. While both image and text augmentations proved effective, the latter performed comparatively better. The reason lies in the high dimensionality and degree of freedom of the image space, which makes it harder to generate augmentations that are both diverse and stay within the distribution. In contrast, the lower-dimensional, structured nature of text makes it more amenable to controlled augmentation. Figure 2 shows that text augmentations can introduce diversity while maintaining a small distribution gap, whereas image augmentations cannot. Additionally, Fig. 3 analyzes the impact of the number of augmentations on multimodal robustness. We observed that increasing text augmentations with I2T(Human) improves robustness, as more captions better approximate the data distribution, however, image augmentations using T-I2I(SD) enhance robustness up to two additional augmentations but can degrade it beyond that due to distribution shift. Therefore, we believe that developing image augmentations with sufficient diversity while minimizing the distribution gap is a promising direction for future research.
Many-to-many augmentations (N:N). Appendix C shows that combining image and text augmentations (N:N) did not enhance performance over text-only augmentations. With two augmented images and four texts, original pairs constitute only a 12.5% of the data, which makes (N:N) augmentations prone to distorting the data distribution if not designed carefully. Thus, while theoretically promising, many-to-many augmentations require a dedicated methodology, which falls out of the scope of this work.
@@
T-I2I(SD)))) 41.1 $\uparrow$0.7 46.0 $\uparrow$1.8 34.8 $\downarrow$0.3
@@ 6.2 Additional tasks
Image captioning (IC). We verify the effectiveness of one-to-many augmentations in IC using the model and dataset originally proposed for this task, BLIP and COCO (Tab. 6). Since IC involves only image inputs, MAT does not surpass FARE, the unimodal AT method tailored for image attacks. Nevertheless, we demonstrate that augmentations enhance robustness in this task, with MAT+ outperforming FARE.
@@ 7 Conclusions
This is the first work to study adversarial defense strategies for vision-language (VL) models against multimodal adversarial attacks in VL tasks. Existing defense methods are ineffective against multimodal attacks, as they focus on defending against image-only attacks, and also ignore the one-to-many nature of images and texts. To address this, we proposed MAT, a novel defense framework that leverages multimodal perturbations and one-to-many augmentations. Our comprehensive analysis reveals that our framework significantly improves adversarial robustness against multimodal attacks and that augmentations are most effective when well-aligned, diverse, and avoid distribution shift; a setting that previous literature on unimodal adversarial defense did not explore. Our novel perspective identified several challenges, such as the difficulty of generating image augmentations without distorting the original joint distribution of images and texts, opening a promising path for future research.
@@ A.1 MAT’s Inner Maximization Strategy
CLIP Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\min_{\delta_{I}}\cos(f_{\theta_{\text{I}}}(I+\delta_{I}),f_{\theta_{\text{T}}}(T))$ $\mathcal{L}_{\text{CLIP}}$
ALBEF Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Visual Grounding $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
BLIP Image-Text Retrieval $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Image Captioning $\max_{\delta_{T}}\|f_{\theta_{T}}(T+\delta_{T})-f_{\theta_{I}}(I)\|$ $\max_{\delta_{I}}\mathcal{L}_{\text{LM}}$ $\mathcal{L}_{\text{LM}}$
@@ A.1.1 Image-Text Retrieval (ITR)
MAT for CLIP generates text adversarial examples by maximizing the divergence between image-text embeddings, using BERT-attack (Li et al., 2020), while image adversarial examples are generated by minimizing their cosine similarity, following image attacks in Co-Attack (Zhang et al., 2022) and SGA attack (Lu et al., 2023).
The ALBEF and BLIP model architectures use a similar approach for ITR but include additional components, such as an ITM module (Image-Text Matching) and multimodal encoders. ITC loss ($\mathcal{L}_{\text{ITC}}$) is the image-text contrastive loss, similar to $\mathcal{L}_{\text{CLIP}}$, with the primary difference being whether momentum encoders are used to store previously seen representations. ITM loss ($\mathcal{L}_{\text{ITM}}$) predicts whether an image-text pair is positive (matched) or negative (not matched), which is predicted by a multimodal encoder in ALBEF and BLIP in addition to their unimodal image and text encoders. Please refer to original papers (Li et al., 2021, 2023) for more details. Given the added complexity in ALBEF and BLIP, MAT for these models is designed as follows:
@@ A.2.2 Visual Grounding (VG)
Dataset. For training, we use MSCOCO’s training set, following the ITR setting. For testing VG tasks, we use RefCOCO+, which includes text descriptions of objects in images along with their corresponding bounding boxes. RefCOCO+ contains 141,564 expressions for 19,992 images from the COCO training set. This dataset was collected interactively through a two-player game (Kazemzadeh et al., 2014). It consists of three test sets, Val, TestA, and TestB. Test Set A contains objects sampled randomly from the entire dataset. Test Set B contains objects sampled from the most frequently occurring object categories in the dataset. Test Set C contains objects sampled from images that contain at least 2 objects of the same category.
Evaluation metric. VG aims to localize the region in an image that corresponds to a specific text description. Following Li et al. (2021), we extend Grad-CAM (Selvaraju et al., 2017) to acquire heatmaps, and use them to rank the detected proposals provided by Yu et al. (2018). We measure the accuracy of the attention maps with the IoU threshold being 0.5.
@@ A.2.3 Image Captioning
Evaluation metric. We use various evaluation metrics to measure the quality of the generated captions, including BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE (Lin & Och, 2004), CIDEr (Vedantam et al., 2015), and SPICE (Anderson et al., 2016).
@@ A.3.1 Text augmentations
LangRW (Language rewrite). Language rewrite (LangRW) (Fan et al., 2024) is a method that rewrites the text data to improve the robustness of the model, using a generative natural language processing model, such as Llama (Touvron et al., 2023). We used Llama-2-7B  [^2] https://huggingface.co/meta-llama/Llama-2-7b. In our work, we used slightly modified prompts from the original work to simultaneously generate four captions per image. Given an original caption $T$, the prompt for generating additional captions are as follows:
     Rewrite image captions in 4 different  ways.    {coco caption 1 for image i}  => {coco caption 2 for image i}  => {coco caption 3 for image i}  => {coco caption 4 for image i}  => {coco caption 5 for image i}    {coco caption 1 for image j}  => {coco caption 2 for image j}  => {coco caption 3 for image j}  => {coco caption 4 for image j}  => {coco caption 5 for image j}    {coco caption 1 for image k}  => {coco caption 2 for image k}  => {coco caption 3 for image k}  => {coco caption 4 for image k}  => {coco caption 5 for image k}    {original caption to be rewritten}  =>     where the coco captions are randomly sampled from the original captions from the COCO dataset (Chen et al., 2015).
I2T(div-Caps) using div-Caps. div-Caps (Chen et al., 2024) is a latest vision-language multimodal model. We use the div-Caps-2.5-2B model. We generate four captions per image using the following four prompts to ensure diversity:
@@
The lower alignment of div-Caps-generated captions compared to original image-text pairs (Fig. 2) is due to the fact that captions generated with Background and Style are less aligned with images than the full captions.
@@ A.3.2 Image augmentations
RandAug (Random Augmentation). RandAug (Cubuk et al., 2020) is an image augmentation method that applies a series of random transformations to the image. We used the codes from ALBEF (Li et al., 2021)  [^3] https://github.com/salesforce/ALBEF. We set the number of operations to 2 and the magnitude to 5 for all experiments.
@@
[…]
@@ A.4 Hyperparameter Details
Appendix B Visualization of Augmentations
Figure 4: Visualization of image augmentations (1).
Figure 5: Visualization of image augmentations (2).
Figure 6: Visualization of text augmentations (1).
Figure 7: Visualization of text augmentations (2).
Appendix C Additional Results
C.1 Image-text retrieval (ITR)
In this section, we present comprehensive results for image-text retrieval (ITR), evaluating defense methods against unimodal attacks (PGD (Madry et al., 2017) for image attack and BERT-attack (Li et al., 2020)) as well as against multimodal attacks (Co-Attack (Zhang et al., 2022) and SGA (Lu et al., 2023)). Table 8 and Table 9 present the results of CLIP trained on Flickr30k and COCO, respectively. Table 10 and Table 11 show the results of ALBEF trained on Flickr30k and COCO, while Table 12 presents the results of BLIP trained on COCO.
@@ C.1 Image-text retrieval (ITR)
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ 6.1.2 Effectiveness of one-to-many augmentations
Through our comprehensive analysis, we reveal that some augmentations consistently enhance robustness, whereas some do not. To identify key factors for effective augmentation, we analyze alignment, diversity, and distribution gap based on our hypothesis for approximating the true image-text distribution (Sec.4.3.2). To analyze this, we introduce three measurable properties:
  • Distribution gap: The Fréchet distance [frechet1936coefficient] between the distributions of the original and augmented samples, $d_{\text{F}}(\mathcal{\hat{D}}_{\text{tr}},\mathcal{\hat{D}}_{\text{aug}})$, where $\mathcal{\hat{D}}_{\text{aug}}={(I_{\text{aug}},T),(I,T_{\text{aug}})}^{N}_{i=1}$. This metric, widely used for evaluating generative models (e.g., FID [heusel2017gans]), quantifies the distribution gap.
Fig. 2 reveals that all properties, high alignment, high diversity, and small distribution gap, is crucial for effective augmentations.
Cross-modal augmentation outperforms intra-modal augmentation due to higher image-text alignment. We observe that cross-modal augmentations yield better robustness than intra-modal ones, by considering the other modality and generating well-aligned image-text pairs. For example, the alignment scores of RandAug and EDA are lower than those of the original pairs (Fig. 2), indicating semantic mismatch (Case. 1). Intra-modal augmentations struggle to balance diversity and alignment: increasing diversity without considering image-text relationships disrupts alignment. Fig. 2 shows that intra-modal augmentations appear in the lower-left, achieving only one of alignment or diversity.
@@
Appendix D provides qualitative examples of the augmentations generated for each technique.
@@
Image augmentations are challenging. While both image and text augmentations proved effective, text performs better. This is because the high-dimensional image space makes it difficult to generate augmentations that are diverse yet distribution-consistent, whereas the structured nature of text allows controlled augmentation. Figure 2 shows that text augmentations can introduce diversity while maintaining a small distribution gap, unlike image augmentations. Additionally, Fig. 3 analyzes the number of augmentations. Increasing text augmentations with I2T(Human) improves robustness, as more captions better approximate the data distribution, however, image augmentations using T-I2I(SD) enhance robustness up to two additional augmentations but can degrade it beyond that due to distribution shift. These results suggest that developing image augmentations with high diversity yet minimal distribution gap is a promising direction for future research.
Many-to-many augmentations (N:N). Appendix B shows that combining image and text augmentations (N:N) did not enhance performance over text-only augmentations. With two augmented images and four texts, original pairs constitute only a 12.5% of the data, which makes (N:N) augmentations prone to distorting the data distribution if not designed carefully. Thus, while theoretically promising, many-to-many augmentations require a dedicated methodology, which falls out of the scope of this work.
@@
T-I2I(SD)) 41.1 $\uparrow$0.7 46.0 $\uparrow$1.8 34.8 $\downarrow$0.3
@@
Ablation study: MAT+ vs. naively increasing data samples. Appendix B.3 (Tab. 14) presents controlled experiments to disentangle the effect of data size from the one-to-many (1:N) strategy. With the same number of data points, MAT+ outperforms the naive 1:1 settings.
@@ 6.2 Additional tasks
Image captioning (IC). We verify the effectiveness of one-to-many augmentations in IC using the model and dataset originally proposed for this task, BLIP and COCO (Tab. 6). Since IC involves only image inputs, MAT does not surpass FARE, the unimodal AT method tailored for image attacks. Nevertheless, augmentations still enhance robustness in this task, with MAT+ outperforming FARE.
@@ 7 Conclusions
This work is the first to study adversarial defense for vision-language (VL) models against multimodal attacks in VL tasks. Existing defenses fail against multimodal attacks as they focus on image-only perturbations and overlook the one-to-many nature of image-text pairs. To address this, we proposed MAT, a novel defense framework that leverages multimodal perturbations and one-to-many augmentations. Through comprehensive analysis, we design efficient yet effective multimodal defenses from an optimization perspective and introduce a data-driven approach to further enhance robustness. Our findings reveal key challenges, such as the difficulty in complex optimization and the need for diverse yet in-distribution augmentations, guiding future research.
Acknowledgment
This work was partially supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grant JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, and by JST K Program Grant JPMJKP24C2 Japan.
@@ A.1 MAT’s Inner Maximization Strategy
CLIP Image-Text Retrieval $\max_{\delta_{T}}\cos(f_{\theta_{T}}(T+\delta_{T}),f_{\theta_{I}}(I))$ $\min_{\delta_{I}}\cos(f_{\theta_{\text{I}}}(I+\delta_{I}),f_{\theta_{\text{T}}}(T))$ $\mathcal{L}_{\text{CLIP}}$
ALBEF Image-Text Retrieval $\max_{\delta_{T}}\cos(f_{\theta_{T}}(T+\delta_{T}),f_{\theta_{I}}(I))$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Visual Grounding $\max_{\delta_{T}}\cos(f_{\theta_{T}}(T+\delta_{T}),f_{\theta_{I}}(I))$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
BLIP Image-Text Retrieval $\max_{\delta_{T}}\cos(f_{\theta_{T}}(T+\delta_{T}),f_{\theta_{I}}(I))$ $\max_{\delta_{I}}\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$ $\mathcal{L}_{\text{ITC}}+\mathcal{L}_{\text{ITM}}$
Image Captioning $\max_{\delta_{T}}\cos(f_{\theta_{T}}(T+\delta_{T}),f_{\theta_{I}}(I))$ $\max_{\delta_{I}}\mathcal{L}_{\text{LM}}$ $\mathcal{L}_{\text{LM}}$
@@ A.1.1 Image-Text Retrieval (ITR)
MAT for CLIP generates text adversarial examples by maximizing the divergence between image-text embeddings, using BERT-attack [li2020bert], while image adversarial examples are generated by minimizing their cosine similarity, following image attacks in Co-Attack [zhang2022towards] and SGA attack [lu2023set].
The ALBEF and BLIP model architectures use a similar approach for ITR but include additional components, such as an ITM module (Image-Text Matching) and multimodal encoders. ITC loss ($\mathcal{L}_{\text{ITC}}$) is the image-text contrastive loss, similar to $\mathcal{L}_{\text{CLIP}}$, with the primary difference being whether momentum encoders are used to store previously seen representations. ITM loss ($\mathcal{L}_{\text{ITM}}$) predicts whether an image-text pair is positive (matched) or negative (not matched), which is predicted by a multimodal encoder in ALBEF and BLIP in addition to their unimodal image and text encoders. Please refer to original papers [li2021align, li2023blip] for more details. Given the added complexity in ALBEF and BLIP, MAT for these models is designed as follows:
@@ A.2.2 Visual Grounding (VG)
Dataset. For training, we use MSCOCO’s training set, following the ITR setting. For testing VG tasks, we use RefCOCO+, which includes text descriptions of objects in images along with their corresponding bounding boxes. RefCOCO+ contains 141,564 expressions for 19,992 images from the COCO training set. This dataset was collected interactively through a two-player game [kazemzadeh2014referitgame]. It consists of three test sets, Val, TestA, and TestB. Test Set A contains objects sampled randomly from the entire dataset. Test Set B contains objects sampled from the most frequently occurring object categories in the dataset. Test Set C contains objects sampled from images that contain at least 2 objects of the same category.
Evaluation metric. VG aims to localize the region in an image that corresponds to a specific text description. Following li2021align, we extend Grad-CAM [selvaraju2017grad] to acquire heatmaps, and use them to rank the detected proposals provided by yu2018mattnet. We measure the accuracy of the attention maps with the IoU threshold being 0.5.
@@ A.2.3 Image Captioning
Evaluation metric. We use various evaluation metrics to measure the quality of the generated captions, including BLEU [papineni2002bleu], METEOR [banerjee2005meteor], ROUGE [lin2004automatic], CIDEr [vedantam2015cider], and SPICE [anderson2016spice].
@@ A.3.1 Text augmentations
LangRW (Language rewrite). Language rewrite (LangRW) [fan2024improving] is a method that rewrites the text data to improve the robustness of the model, using a generative natural language processing model, such as Llama [touvron2023llama]. We used Llama-2-7B  [^2] https://huggingface.co/meta-llama/Llama-2-7b. In our work, we used slightly modified prompts from the original work to simultaneously generate four captions per image. Given an original caption $T$, the prompt for generating additional captions are as follows:
      Rewrite image captions in 4 different    ways.        {coco caption 1 for image i}    => {coco caption 2 for image i}    => {coco caption 3 for image i}    => {coco caption 4 for image i}    => {coco caption 5 for image i}        {coco caption 1 for image j}    => {coco caption 2 for image j}    => {coco caption 3 for image j}    => {coco caption 4 for image j}    => {coco caption 5 for image j}        {coco caption 1 for image k}    => {coco caption 2 for image k}    => {coco caption 3 for image k}    => {coco caption 4 for image k}    => {coco caption 5 for image k}        {original caption to be rewritten}    =>      where the coco captions are randomly sampled from the original captions from the COCO dataset [chen2015microsoft].
I2T(div-Caps) using InternVL. InternVL [chen2024internvl] is a latest vision-language multimodal model. We use the InternVL-2.5-2B model. We generate four captions per image using the following four prompts to ensure diversity:
@@
The lower alignment of InternVL-generated captions compared to original image-text pairs (Fig. 2) is due to the fact that captions generated with Background and Style are less aligned with images than the full captions.
@@ A.3.2 Image augmentations
RandAug (Random Augmentation). RandAug [cubuk2020randaugment] is an image augmentation method that applies a series of random transformations to the image. We used the codes from ALBEF [li2021align]  [^3] https://github.com/salesforce/ALBEF. We set the number of operations to 2 and the magnitude to 5 for all experiments.
@@
Stable Diffusion (SD); Text-guided Image-to-Image. We used SD-v2.1, and use the pipeline for text-guided image-to-image generation [meng2021sdedit] from Huggingface [^5] https://huggingface.co/docs/diffusers/api/pipelines/stable_diffusion/img2img. We used a strength hyperparameter of 0.5, which controls the diversity of augmentations. Larger strength increases diversity, however, caused a distribution shift that negatively impacted performance.
@@ A.4 Hyperparameter Details
A.5 Computational Cost
All experiments were conducted on a single NVIDIA A100 GPU.
[…]
In MAT+, the cost of the augmentations depends on each technique. Cross-modal augmentations using generative models (e.g., InternVL for captioning, Stable Diffusion for image-to-image generation) require more time, taking approximately 3 days on COCO. However, the augmentations are created in advance only once, so their generation has no effect on the actual cost of the adversarial training.
Appendix B Additional Results
B.1 Image-text retrieval (ITR)
In this section, we present comprehensive results for image-text retrieval (ITR), evaluating defense methods against unimodal attacks (PGD [madry2017towards] for image attack and BERT-attack [li2020bert]) as well as against multimodal attacks (Co-Attack [zhang2022towards] and SGA [lu2023set]). Table 8 and Table 9 present the results of CLIP trained on Flickr30k and COCO, respectively. Table 10 and Table 11 show the results of ALBEF trained on Flickr30k and COCO, while Table 12 presents the results of BLIP trained on COCO.
@@ B.1 Image-text retrieval (ITR)
(ours) MAT+ Basic(EDA) 85.4 $\uparrow$1.7 69.5 $\uparrow$2.0 79.5 $\uparrow$2.1 62.4 $\uparrow$1.0 75.7 $\uparrow$3.5 53.7 $\uparrow$2.6 60.1 $\uparrow$3.2 39.5 $\uparrow$2.4 39.1 $\uparrow$1.6 27.5 $\uparrow$2.7
@@
(ours) MAT+ Basic(EDA) 55.9 $\uparrow$0.2 40.2 $\downarrow$0.5 48.6 $\downarrow$0.9 35.3 $\downarrow$0.8 41.1 $\downarrow$1.8 27.1 $\downarrow$0.4 29.8 $\downarrow$1.6 19.5 $\downarrow$0.0 18.4 $\uparrow$0.7 12.9 $\uparrow$0.6
@@
(ours) MAT+ Basic(EDA) 82.2 $\uparrow$0.2 67.9 $\uparrow$1.6 78.2 $\uparrow$0.2 63.8 $\uparrow$0.0 78.6 $\uparrow$1.5 58.1 $\uparrow$2.0 71.1 $\downarrow$1.9 52.9 $\downarrow$1.1 44.6 $\downarrow$2.5 31.2 $\downarrow$1.7
@@
(ours) MAT+ Basic(EDA) 63.9 $\uparrow$0.0 46.8 $\uparrow$0.6 53.2 $\downarrow$7.0 38.7 $\uparrow$0.4 54.4 $\downarrow$1.0 36.4 $\downarrow$0.5 50.8 $\downarrow$2.2 30.9 $\downarrow$1.2 31.5 $\uparrow$0.3 20.9 $\downarrow$0.3
@@
(ours) MAT+ I2T(Human) 71.0 $\uparrow$4.1 54.3 $\uparrow$4.5 68.6 $\uparrow$0.8 52.1 $\uparrow$2.6 57.1 $\uparrow$2.4 40.4 $\uparrow$3.3 51.2 $\uparrow$0.7 35.3 $\uparrow$2.5 35.6 $\uparrow$4.3 25.7 $\uparrow$4.7